# The impact of assembly history on the X-ray detectability of halos

### v1

@@
I. Marini ilaria.marini@eso.org11    P. Popesso 11 2 2    K. Dolag 33 4 4 2 2    V. Biffi 55 6 6    S. Vladutescu-Zopp 44    T. Castro 55 6 6 7 7 8 8    V. Toptun 11    N. de Isídio 11    A. Dev $1$    D. Mazengo 11 12 12    J. Comparat 1313    C. Gouin 99 10 10    N. Malavasi 1313    A. Merloni 1313    T. Mroczkowski 11
@@ G. Ponti
(Received - ; accepted - )
Galaxy groups represent a significant fraction of the halo population, playing a crucial role in galaxy formation and evolution. However, their detection in X-rays remains challenging, raising questions about the physical mechanisms driving their detectability in current surveys.
Using the Magneticum simulations, we construct a mock X-ray lightcone of the local Universe ($z<0.2$) to investigate the selection function of galaxy groups and clusters. We find that AGN activity is a key driver of baryon depletion, but late-time mergers boost X-ray brightness by replenishing the gas reservoir in the halos, highlighting the interplay between feedback processes and the environment. Our analysis shows that X-ray bright groups experience sustained late-time mass accretion, maintaining higher gas fractions and fuelling the central supermassive black holes (SMBH), further increasing the X-ray emissivity in the core. In contrast, X-ray faint groups form earlier and lose most of their gas over time, resembling fossil groups.
Magneticum predicts strong anti-correlations between gas fraction (or X-ray luminosity) and SMBH mass, stellar mass (both in the central galaxy and intracluster light), and group richness at fixed halo mass. We derive predictions on the hot gas fraction at fixed halos mass (e.g. a group of total mass $M_{500}=10^{13}M_{\odot}$ can have hot gas fractions in the range $f_{\mathrm{gas}}=0.02-0.06$ and a central SMBH with a median mass of $M_{\mathrm{BH}}=10^{9}M_{\odot}$ and a scatter of $0.5$ dex) compatible with the most recent measurements of the baryonic fraction. These findings will aid the interpretation of future X-ray surveys, demonstrating the power of simulation-based inference.
@@
Figure 1: A schematic representation of the X-ray lightcone LC30, where we show the distribution of galaxy groups and clusters with $M_{200}>10^{12.5}M_{\odot}$ and colour-coded for their intrinsic X-ray luminosity in $0.5-2.0$ keV within $R_{500}$. The X-ray emission is based on PHOX’s photon list. Each smaller panel represents the projected X-ray emission corresponding to the shells at different snapshots forming LC30 and has an angular scale of $30\times 30$ square degrees.
@@ 1 Introduction
One of the primary ways to study galaxy groups is through their X-ray emission, which originates from hot, diffuse gas in the IGrM. Moreover, at low temperatures, line cooling becomes very important, and the emissivity (both in soft and bolometric bands) becomes strongly metallicity- and temperature-dependent (see Fig. 1 in Lovisari et al. 2021). In this context, X-ray detectability refers to the ability to identify galaxy groups through their extended X-ray emission, which depends on factors such as gas density, temperature, metallicity, and the sensitivity of current and future X-ray observatories. Observations from missions such as Chandra, XMM-Newton, and eROSITA have revealed significant variations in the X-ray properties of groups, with some systems appearing bright and extended (Sun et al. 2009; Andreon et al. 2017; Bulbul et al. 2024; Bahar et al. 2024). In contrast, many other galaxy groups remain faint or undetected (Zheng et al. 2023; Popesso et al. 2024a, c; Marini et al. 2024, 2025b; Li et al. 2024). While an optical follow-up to detect them is possible (Bulbul et al. 2022; Ota et al. 2023; Balzer et al. 2025), a considerable theoretical effort is in place to understand such systems (e.g. Oppenheimer et al. 2020; Ragagnin et al. 2022; Cui et al. 2024; Hough et al. 2024; Bigwood et al. 2025). These differences are thought to be due to the structural properties of groups, driven by non-gravitational processes that alter the thermodynamic state of the hot gas and the overall baryon content in groups and clusters (Wu et al. 2000; Borgani et al. 2002; Davies et al. 2019; Oppenheimer et al. 2020). In principle, their assembly history, merger activity, and impact of feedback from active galactic nuclei (AGN) could all affect the hot gas distribution, while stellar feedback is negligible (e.g. Bigwood et al. 2025). However, the key mechanisms driving the emergence of X-ray bright and faint populations remain an open question. This paper addresses this issue by exploring several processes potentially responsible for these distinct populations, providing new insights into their formation and evolution.
@@
AGN feedback plays a central role in shaping the galaxy luminosity function and regulating the thermal and dynamical properties of halos (e.g. Planelles et al. 2014). Supermassive black holes (SMBHs) at the centres of galaxies can inflate large energetic bubbles that couple with the ICM and IGrM, pushing the inner gas to great distances. Additionally, outflows and jets inject vast amounts of energy into the IGrM, preventing gas reservoirs from cooling efficiently and suppressing star formation (McNamara & Nulsen 2007; Gaspari et al. 2020). Shock fronts distribute this heat throughout the environment, and in galaxy groups, the total feedback energy may become comparable to or even exceed the halo’s binding energy (see an extensive review Eckert et al. 2021). While in galaxy clusters, this effect invests only the core region, in low-mass halos it extends to the virial scale (e.g. Sorini et al. 2022; Angelinelli et al. 2022, 2023; Ayromlou et al. 2023). State-of-the-art galaxy evolution models universally incorporate AGN feedback. Yet different implementations can yield drastically different predictions (Habouzit et al. 2021; Schaye et al. 2023), varying from exceedingly hot gas-rich groups to systems devoid of gas (with a direct effect on the X-ray appearance of these systems on a global scale, see Popesso et al. 2024a).
@@
In this work, we leverage the Magneticum Pathfinder simulations (Hirschmann et al. 2014; Dolag 2015, Dolag in prep.) a suite of state-of-the-art cosmological hydrodynamical simulations, to investigate the physical mechanisms behind the X-ray selection of galaxy groups. Magneticum provides a powerful framework for exploring how different assembly histories (i.e. halo accretion history) and feedback mechanisms lead to variations in gas content, BH growth, and X-ray luminosity. Unlike other cosmological simulations (e.g. SIMBA, Eagle, and IllustrisTNG; see Habouzit et al. 2021), Magneticum reproduces the AGN luminosity function well down to $z=2$ (Hirschmann et al. 2014; Steinborn et al. 2018; Biffi et al. 2018), which is responsible for most of the X-ray cosmic background, and recovers several scaling relations obtained with eROSITA at the group and galaxy mass scales (Bahar et al. 2022; Zhang et al. 2024b). Magneticum achieves this without an explicit calibration on observed scaling relations down to the group-scale halos, in contrast to other state-of-the-art simulations, such as the IllustrisTNG project (Pillepich et al. 2018; Nelson et al. 2018), SIMBA (Davé et al. 2019), and FLAMINGO (Schaye et al. 2023). Thus, we argue that the Magneticum suite is an excellent tool to infer predictions of the hot gas in galaxy groups and investigate the detectability of such objects. By analysing synthetic X-ray observations and tracking the evolution of key properties such as gas fraction, environment, and SMBH activity, we aim to uncover the driving forces behind the X-ray selection of groups. Our findings may have important implications for the interpretation of X-ray surveys, the study of galaxy evolution, and the development of future observational strategies.
@@ 2 Simulations
This analysis is based on synthetic data from the X-ray lightcone described in Marini et al. (2024, 2025b, 2025a) and the original Box2/hr at $z=0$ of the Magneticum set. The Magneticum Pathfinder simulation [^1] http://www.magneticum.org/index.html is an extended set of the state-of-the-art cosmological hydrodynamical simulations carried out with P-GADGET3 (an updated version of the public GADGET-2 code Springel 2005). Key advancements include a higher-order kernel function, time-dependent artificial viscosity and artificial conduction schemes (Dolag et al. 2005; Beck et al. 2016).
@@ 2.2 The LC30 lightcone and X-ray emission
Complete geometric and physical description of the lightcone is provided in Marini et al. (2024, 2025b, 2025a), here we briefly summarise the main points. Figure 1 illustrates the distribution of groups and clusters (with halo mass [^2] We define $M_{\Delta}$ as the mass encompassed by a mean overdensity equal to $\Delta$ times the critical density of the universe $\rho_{c}(z)$. $M_{200}>10^{12.5}M_{\odot}$) colour-coded for their X-ray luminosity within $R_{500}$. The lightcone spans $30\times 30$ square degrees (LC30) and consists of five sub-cubes extracted from the most recent five snapshots in Box2/hr. The projected distribution of photons ($0.5-2.0$ keV) of each sub-cube is illustrated in the subpanels of Fig. 1. Halos and substructures in the simulations are identified in post-processing using a two-step approach: first, a Friend-of-Friend (FOF) algorithm is applied, followed by SubFind (Springel et al. 2001; Dolag et al. 2009) to catalogue the subhalos. The X-ray emission is modelled using PHOX (Biffi et al. 2012, 2013, 2018; Vladutescu-Zopp et al. 2023) which simulates X-ray spectral emission for the gas particles, the BH, and the star particles (i.e. X-ray binaries) using spectral models from the XSPEC library (v12; Arnaud 1996). A fiducial collecting area $A_{fid}$ and exposure time $\tau_{fid}$ are initially assumed to generate a statistically significant number of photons from the spectra using a Monte Carlo approach. These photons are then blueshifted or redshifted based on their location and projection within the lightcone. For this analysis, we do not use the eROSITA mock observation (similarly to Marini et al. 2024, 2025b), instead we calculate the X-ray luminosities from the theoretical PHOX’s photon lists. These are estimated in the rest-frame $0.5-2.0$ keV energy range.
@@ 3 Explaining the X-ray selection in galaxy groups
Despite the paramount importance of galaxy groups from the astrophysical and cosmological point of view, the challenges in their detection have limited their studies. Many studies have been devoted to clarifying the X-ray selection function of X-ray surveys like eROSITA (Clerc et al. 2018; Liu et al. 2022; Seppi et al. 2022; Clerc et al. 2024; Marini et al. 2024, 2025b). Here, we follow up on the work in Marini et al. (2024, 2025b) which showed the selection function in X-ray surveys at the nominal depth of eRASS:4, highlighting a primary bias against galaxy groups with lower surface brightness and high core entropy. The authors also examined the scatter in the $L_{500}-M_{500}$ relation within Magneticum, calculating the residuals $\log\Delta L_{500}$ from the Lovisari et al. (2015) $L_{500}-M_{500}$ scaling relation – namely $\log\Delta L_{500}$$\equiv\log(L_{500}/L_{500,\,BF})$, where $L_{500}$ is the intrinsic luminosity and $L_{500,\,BF}$ is the best-fit value – given that it provides the smallest residuals among the literature scaling relations compared in LC30 (Marini et al. 2024, 2025b). The simulation indicates that the groups in the upper envelope of the relation exhibit a more dense gas distribution in the core (i.e. $0.15R_{500}$) than those in the lower envelope within a fixed halo mass bin.
@@
This section investigates whether the central hot gas [^3] Here, and throughout the rest of the paper, we refer to hot gas as all the non-star-forming gas particles in the simulations with temperatures above $5\times 10^{5}K$. Additionally, we use the subscript “$f_{\mathrm{gas}}$” to refer to hot gas and “$f_{\mathrm{gas,\,tot}}$” for both hot and cold gas. concentration is the sole factor potentially enhancing or diminishing the total X-ray emission in galaxy groups and clusters in Magneticum. Under this hypothesis, the central region (e.g. within $0.15R_{500}$) should account for most of the total X-ray emission and should define cool core systems (González Villalba et al. 2025). We test this in Fig 2, where we illustrate the ratio between the central and total X-ray luminosity from the IGrM only (no XRB or AGN) of individual groups and clusters within LC30 at fixed halo mass as a function of $\log\Delta L_{500}$. To account for the halo mass dependence in the $L_{500}-M_{500}$ relation, we maintain the $\log\Delta L_{500}$ definition which measures the distance between the halo’s measured X-ray luminosity $L_{500}$ (in rest-frame $0.5-2.0$ keV) from the best-fit power-law scaling relation ($L_{500,\,BF}$) given in Lovisari et al. (2015) scaling relation (see Marini et al. 2024, 2025b). Values of $\log\Delta L_{500}$ larger than zero correspond to over-luminous galaxy groups and vice-versa. Additionally, we define a luminosity concentration as the ratio between the luminosity calculated within $0.15R_{500}$ and $R_{500}$. In Fig. 2, we quantify the importance of the central BH activity (left panel) and the hot gas content in $R_{500}$ (right panel) in shaping this distribution for objects with $M_{500}>10^{12.5}M_{\odot}$. The colour bar encompasses the 5–95th percentiles. To guide the eye, we overplot two dotted black lines over the $\log\Delta L_{500}$$=0$ and $L(<0.15R_{500})/L_{500}=0.5$ relations.
@@
Figure 5: Evolution of several halo properties for a sample of bright and faint halos in the halo mass range $M_{500}=10^{13}-10^{13.5}M_{\odot}$ at $z=0$. Left panel: Median baryonic fraction (solid lines) across different redshifts. We also report the hot gas contribution (dashed) and total gas (dashed-dotted). The fractions are calculated within $R_{500}$ of the group at the corresponding redshift. Right panel: Median halo mass accretion as a function of the simulation redshift. The dashed-dotted lines represent the contribution only from the DM. We also report the pseudo-evolution of a static NFW profile in the range of concentration of our simulated halos $103<c_{500}<$ 10. The legend is common in all panels: the halo samples are selected from the original Magneticum cosmological box (i.e. Box2/hr). The shaded regions mark the 16–84th percentile region.

### v2

@@
I. Marini Corresponding author: ilaria.marini@eso.org11    P. Popesso 11 2 2    K. Dolag 33 4 4 2 2    V. Biffi 55 6 6    S. Vladutescu-Zopp 33    T. Castro 55 6 6 7 7 8 8    V. Toptun 11    N. de Isídio 11    A. Dev $1$    D. Mazengo 11 12 12    J. Comparat 1313    C. Gouin 99 10 10    N. Malavasi 1313    A. Merloni 1313    T. Mroczkowski 11
@@ G. Ponti
(Received 21 March 2025 ; accepted 17 April 2025 )
Galaxy groups represent a significant fraction of the halo population, playing a crucial role in galaxy formation and evolution. However, their detection in X-rays remains challenging, raising questions about the physical mechanisms driving their detectability in current surveys. Using the Magneticum simulations, we construct a mock X-ray lightcone of the local Universe ($z<0.2$) to investigate the selection function of galaxy groups and clusters. We find that the central supermassive black holes (SMBH) activity is a key driver of baryon depletion, but late-time mergers boost X-ray brightness by replenishing the gas reservoir in the halos, highlighting the interplay between feedback processes and the environment. Our analysis shows that X-ray bright groups experience sustained late-time mass accretion, maintaining higher gas fractions and fuelling the central SMBH, further increasing the X-ray emissivity in the core. In contrast, X-ray faint groups form earlier and lose most of their gas over time, resembling fossil groups. These faint groups reside preferentially in under-dense regions, unable to re-accrete efficiently baryons at later times. Magneticum predicts strong anti-correlations between gas fraction (or X-ray luminosity) and SMBH mass, stellar mass (both in the central galaxy and intracluster light), and group richness at fixed halo mass. We derive predictions on the hot gas fraction at fixed halos mass (e.g. a group of total mass $M_{500}=10^{13}M_{\odot}$ can have hot gas fractions in the range $f_{\mathrm{gas}}=0.02-0.06$ and a central SMBH with a median mass of $M_{\mathrm{BH}}=10^{9}M_{\odot}$ and a scatter of $0.5$ dex) compatible with the most recent measurements of the baryonic fraction. These findings will aid the interpretation of future X-ray surveys, demonstrating the power of simulation-based inference.
@@
Figure 1: Schematic representation of the X-ray lightcone LC30, where we show the distribution of galaxy groups and clusters with $M_{200}>10^{12.5}M_{\odot}$ and colour-coded for their intrinsic X-ray luminosity in $0.5-2.0$ keV within $R_{500}$. The X-ray emission is based on PHOX’s photon list. Each smaller panel represents the projected X-ray emission corresponding to the shells at different snapshots forming LC30 and has an angular scale of $30\times 30$ square degrees.
@@ 1 Introduction
One of the primary ways to study galaxy groups is through their X-ray emission, which originates from hot, diffuse gas in the IGrM. Moreover, at low temperatures, line cooling becomes very important, and the emissivity (both in soft and bolometric bands) becomes strongly metallicity- and temperature-dependent (see Fig. 1 in Lovisari et al. 2021). In this context, X-ray detectability refers to the ability to identify galaxy groups through their extended X-ray emission, which depends on factors such as gas density, temperature, metallicity, and the sensitivity of current and future X-ray observatories. Observations from missions such as Chandra, XMM-Newton, and eROSITA have revealed significant variations in the X-ray properties of groups, with some systems appearing bright and extended (Sun et al. 2009; Andreon et al. 2017; Bulbul et al. 2024; Bahar et al. 2024). In contrast, many other galaxy groups remain faint or undetected (Zheng et al. 2023; Popesso et al. 2024a, c; Marini et al. 2024, 2025; Li et al. 2024). While an optical follow-up to detect them is possible (Bulbul et al. 2022; Ota et al. 2023; Balzer et al. 2025), a considerable theoretical effort is in place to understand such systems (e.g. Oppenheimer et al. 2020; Ragagnin et al. 2022; Cui et al. 2024; Hough et al. 2024; Bigwood et al. 2025). These differences are thought to be due to the structural properties of groups, driven by non-gravitational processes that alter the thermodynamic state of the hot gas and the overall baryon content in groups and clusters (Wu et al. 2000; Borgani et al. 2002; Davies et al. 2019; Oppenheimer et al. 2020). In principle, their assembly history, merger activity, and impact of feedback from active galactic nuclei (AGN) could all affect the hot gas distribution, while stellar feedback is negligible (e.g. Bigwood et al. 2025). However, the key mechanisms driving the emergence of X-ray bright and faint populations remain an open question. This paper addresses this issue by exploring several processes potentially responsible for these distinct populations, providing new insights into their formation and evolution.
@@
The AGN feedback plays a central role in shaping the galaxy luminosity function and regulating the thermal and dynamical properties of halos (e.g. Planelles et al. 2014). Supermassive black holes (SMBHs) at the centres of galaxies can inflate large energetic bubbles that couple with the ICM and IGrM, pushing the inner gas to great distances. Additionally, outflows and jets inject vast amounts of energy into the IGrM, preventing gas reservoirs from cooling efficiently and suppressing star formation (McNamara & Nulsen 2007; Gaspari et al. 2020). Shock fronts distribute this heat throughout the environment, and in galaxy groups, the total feedback energy may become comparable to or even exceed the halo’s binding energy (see an extensive review Eckert et al. 2021). While in galaxy clusters, this effect invests only the core region, in low-mass halos it extends to the virial scale (e.g. Sorini et al. 2022; Angelinelli et al. 2022, 2023; Ayromlou et al. 2023). State-of-the-art galaxy evolution models universally incorporate AGN feedback. Yet different implementations can yield drastically different predictions (Habouzit et al. 2021; Schaye et al. 2023), varying from exceedingly hot gas-rich groups to systems devoid of gas (with a direct effect on the X-ray appearance of these systems on a global scale, see Popesso et al. 2024a).
@@
In this work, we leverage the Magneticum Pathfinder simulations (Hirschmann et al. 2014; Dolag et al. 2025), a suite of state-of-the-art cosmological hydrodynamical simulations, to investigate the physical mechanisms behind the X-ray selection of galaxy groups. Magneticum provides a powerful framework for exploring how different assembly histories (i.e. halo accretion history) and feedback mechanisms lead to variations in gas content, BH growth, and X-ray luminosity. Unlike other cosmological simulations (e.g. SIMBA, Eagle, and IllustrisTNG; see Habouzit et al. 2021), Magneticum reproduces the AGN luminosity function well down to $z=2$ (Hirschmann et al. 2014; Steinborn et al. 2018; Biffi et al. 2018), which is responsible for most of the X-ray cosmic background, and recovers several scaling relations obtained with eROSITA at the group and galaxy mass scales (Bahar et al. 2022; Zhang et al. 2024b). Magneticum achieves this without an explicit calibration on observed scaling relations down to the group-scale halos, in contrast to other state-of-the-art simulations, such as the IllustrisTNG project (Pillepich et al. 2018; Nelson et al. 2018), SIMBA (Davé et al. 2019), and FLAMINGO (Schaye et al. 2023). Thus, we argue that the Magneticum suite is an excellent tool to infer predictions of the hot gas in galaxy groups and investigate the detectability of such objects. By analysing synthetic X-ray observations and tracking the evolution of key properties such as gas fraction, environment, and SMBH activity, we aim to uncover the driving forces behind the X-ray selection of groups. Our findings may have important implications for the interpretation of X-ray surveys, the study of galaxy evolution, and the development of future observational strategies.
@@ 2 Simulations
This analysis is based on synthetic data from the X-ray lightcone described in Marini et al. (2024, 2025, 2025) and the original Box2/hr at $z=0$ of the Magneticum set. The Magneticum Pathfinder simulation [^1] http://www.magneticum.org/index.html is an extended set of the state-of-the-art cosmological hydrodynamical simulations carried out with P-GADGET3 (an updated version of the public GADGET-2 code Springel 2005). Key advancements include a higher-order kernel function, time-dependent artificial viscosity and artificial conduction schemes (Dolag et al. 2005; Beck et al. 2016).
@@ 2.2 The LC30 lightcone and X-ray emission
Complete geometric and physical description of the lightcone is provided in Marini et al. (2024, 2025, 2025), here we briefly summarise the main points. Figure 1 illustrates the distribution of groups and clusters (with halo mass [^2] We define $M_{\Delta}$ as the mass encompassed by a mean overdensity equal to $\Delta$ times the critical density of the universe $\rho_{c}(z)$. $M_{200}>10^{12.5}M_{\odot}$) colour-coded for their X-ray luminosity within $R_{500}$. The lightcone spans $30\times 30$ square degrees (LC30) and consists of five sub-cubes extracted from the most recent five snapshots in Box2/hr. The projected distribution of photons ($0.5-2.0$ keV) of each sub-cube is illustrated in the subpanels of Fig. 1. Halos and substructures in the simulations are identified in post-processing using a two-step approach: first, a Friend-of-Friend (FOF) algorithm is applied, followed by SubFind (Springel et al. 2001; Dolag et al. 2009) to catalogue the subhalos. The X-ray emission is modelled using PHOX (Biffi et al. 2012, 2013, 2018; Vladutescu-Zopp et al. 2023) which simulates X-ray spectral emission for the gas particles, the BH, and the star particles (i.e. X-ray binaries) using spectral models from the XSPEC library (v12; Arnaud 1996). A fiducial collecting area $A_{fid}$ and exposure time $\tau_{fid}$ are initially assumed to generate a statistically significant number of photons from the spectra using a Monte Carlo approach. These photons are then blueshifted or redshifted based on their location and projection within the lightcone. For this analysis, we do not use the eROSITA mock observation (similarly to Marini et al. 2024, 2025), instead we calculate the X-ray luminosities from the theoretical PHOX’s photon lists. These are estimated in the rest-frame $0.5-2.0$ keV energy range.
@@ 3 Explaining the X-ray selection in galaxy groups
Despite the paramount importance of galaxy groups from the astrophysical and cosmological point of view, the challenges in their detection have limited their studies. Many studies have been devoted to clarifying the X-ray selection function of X-ray surveys like eROSITA (Clerc et al. 2018; Liu et al. 2022; Seppi et al. 2022; Clerc et al. 2024; Marini et al. 2024, 2025). Here, we follow up on the work in Marini et al. (2024, 2025) which showed the selection function in X-ray surveys at the nominal depth of eRASS:4, highlighting a primary bias against galaxy groups with lower surface brightness and high core entropy. The authors also examined the scatter in the $L_{500}-M_{500}$ relation within Magneticum, calculating the residuals $\log\Delta L_{500}$ from the Lovisari et al. (2015) $L_{500}-M_{500}$ scaling relation – namely $\log\Delta L_{500}$$\equiv\log(L_{500}/L_{500,\,BF})$, where $L_{500}$ is the intrinsic luminosity and $L_{500,\,BF}$ is the best-fit value – given that it provides the smallest residuals among the literature scaling relations compared in LC30 (Marini et al. 2024, 2025). The simulation indicates that the groups in the upper envelope of the relation exhibit a more dense gas distribution in the core (i.e. $0.15R_{500}$) than those in the lower envelope within a fixed halo mass bin.
@@
This section investigates whether the central hot gas [^3] Here, and throughout the rest of the paper, we refer to hot gas as all the non-star-forming gas particles in the simulations with temperatures above $5\times 10^{5}K$. Additionally, we use the subscript “$f_{\mathrm{gas}}$” to refer to hot gas and “$f_{\mathrm{gas,\,tot}}$” for both hot and cold gas. concentration is the sole factor potentially enhancing or diminishing the total X-ray emission in galaxy groups and clusters in Magneticum. Under this hypothesis, the central region (e.g. within $0.15R_{500}$) should account for most of the total X-ray emission and should define cool core systems (González Villalba et al. 2025). We test this in Fig 2, where we illustrate the ratio between the central and total X-ray luminosity from the IGrM only (no XRB or AGN) of individual groups and clusters within LC30 at fixed halo mass as a function of $\log\Delta L_{500}$. To account for the halo mass dependence in the $L_{500}-M_{500}$ relation, we maintain the $\log\Delta L_{500}$ definition which measures the distance between the halo’s measured X-ray luminosity $L_{500}$ (in rest-frame $0.5-2.0$ keV) from the best-fit power-law scaling relation ($L_{500,\,BF}$) given in Lovisari et al. (2015) scaling relation (see Marini et al. 2024, 2025). Values of $\log\Delta L_{500}$ larger than zero correspond to over-luminous galaxy groups and vice-versa. Additionally, we define a luminosity concentration as the ratio between the luminosity calculated within $0.15R_{500}$ and $R_{500}$. In Fig. 2, we quantify the importance of the central BH activity (left panel) and the hot gas content in $R_{500}$ (right panel) in shaping this distribution for objects with $M_{500}>10^{12.5}M_{\odot}$. The colour bar encompasses the 5–95th percentiles. To guide the eye, we overplot two dotted black lines over the $\log\Delta L_{500}$$=0$ and $L(<0.15R_{500})/L_{500}=0.5$ relations.
@@
Figure 5: Evolution of several halo properties for a sample of bright and faint halos in the halo mass range $M_{500}=10^{13}-10^{13.5}M_{\odot}$ at $z=0$. Left panel: Median baryonic fraction (solid lines) across different redshifts. We also report the hot gas contribution (dashed) and total gas (dashed-dotted). The fractions are calculated within $R_{500}$ of the group at the corresponding redshift. Right panel: Median halo mass as a function of the simulation redshift. The dashed-dotted lines represent the contribution only from the DM. We also report the pseudo-evolution of a static NFW profile in the range of concentration of our simulated halos $103<c_{500}<$ 10. The legend is common in all panels: the halo samples are selected from the original Magneticum cosmological box (i.e. Box2/hr). The shaded regions mark the 16–84th percentile region.